Charles has a mortgage on his house, with principal $103,500, at 6.75% per annum amortized over 25 years. The house is assessed at $129,500, and the mill rate is 16.3081. Calculate and list his monthly housing costs.
Answer by ewatrrr(24785)


Hi, previously posted
M = P [ i(1 + i)^n ] / [ (1 + i)^n - 1]
Note: i = .0675/12 = .005625 and (1+i)^n = (1.005625)^300 = 5.38
M = 103500*.005625*5.38/ 4.38 = 715.10, is the mortagage payment
Propery Tax = 16.3081*129500/1000 = 2111.90/yr OR 175.99/mo
his monthly housing costs = M + PTax = 715.10 + 175.99